# xT: Nested Tokenization for Larger Context in Large Images

### v1

@@ Abstract
Modern computer vision pipelines handle large images in one of two sub-optimal ways: down-sampling or cropping. These two methods incur significant losses in the amount of information and context present in an image. There are many downstream applications in which global context matters as much as high frequency details, such as in real-world satellite imagery; in such cases researchers have to make the uncomfortable choice of which information to discard. We introduce xT, a simple framework for vision transformers which effectively aggregates global context with local details and can model large images end-to-end on contemporary GPUs. We select a set of benchmark datasets across classic vision tasks which accurately reflect a vision model’s ability to understand truly large images and incorporate fine details over large scales and assess our method’s improvement on them. By introducing a nested tokenization scheme for large images in conjunction with long-sequence length models normally used for natural language processing, we are able to increase accuracy by up to 8.6% on challenging classification tasks and $F_{1}$ score by 11.6 on context-dependent segmentation in large images.
@@ 1 Introduction
Images have been getting increasingly larger over the past decade. Images captured by consumer cameras on smartphones now capture images at 4K resolution (roughly 8.3M pixels) while professional DSLR cameras capture images at 8K resolution. Elsewhere, sensors on satellites and microscopes capture images with over a billion pixels.
@@
We introduce xT, a framework by which myopic vision backbones can effectively integrate local and global context over large images. In particular,we tackle both issues of quadratically-increasing GPU memory utilization and the integration of context across very large images. We achieve this by improving the strengths of hierarchical vision backbones [21, 30] through nested tokenization of images and processing resulting features with long-sequence models, such as Transformer-XL and Mamba [3, 9], from the field of natural language processing. xT matches and beats the performance of competitive large image architectures on multiple downstream tasks that require large visual contexts such as segmentation, detection, and classification. We demonstrate results on a variety of downstream tasks and achieve up to an 8.6% gain in accuracy on classification tasks and an 11.6 increase in $F_{1}$ score on context-dependent segmentation.
@@ 2 Related Works
We evaluate on real-world datasets such as xView3-SAR [25], a dataset where the average image is $29,400\times 24,400$ pixels large, and iNaturalist 2018 [34], where the entire $800\times 800$ image is utilized at once for evaluation. We also evaluate on the standard MS-COCO dataset for direct comparison to vision backbones built for smaller images.
@@
xT extracts features from windows across a large image iteratively using a vision backbone. These features are fed into a shallow context encoder which integrates local context globally, effectively increasing the receptive field of the vision backbone across the entire image while staying within memory and parameter limits.
@@ 4 Methodology
Our goal for xT is to demonstrate a simple framework for allowing existing methods to process large images in a memory efficient and contextual manner. We achieve this through an iterative, two-stage design. First, images are tokenized hierarchically (Section 4.1) before being independently featurized by a region encoder with a limited context window (Section 4.2). Then, a lightweight context encoder incorporates context globally across this sequence of features (Section 4.3), which then gets passed to the task-specific decoders. Our overall pipeline is illustrated in Figure 2.
@@ 4.2 Region Encoder
We start with an image model which is trained for a typical size of small images $H\times W$, usually $224\times 224$ or $256\times 256$. The region encoder independently generates feature maps for each region $R^{i}_{1...P}$. Our region encoders are hierarchical vision transformers in which the sequence length decreases from the beginning of the process, and is also less than the equivalent length produced by isotropic ViTs [5]. We are therefore able to combine more region features together than we would with other architectures, as our sequence length is reduced by 4×4$\times$4 × or greater.
When GPU memory allows, we compute features for all regions simultaneously. However, when the image is too large such that all of its constituent regions cannot fit into GPU memory, we process each region sequentially.
@@ 4.3 Context Encoder
The output region features are concatenated in row major order to form a sequence. In the xT framework, the context encoder is a lightweight sequence-to-sequence model that attends across long sequence lengths. Critically, we constrain the context encoder to have a near-linear memory and time cost. This design allows xT to see many other regions of the large image which otherwise would not be feasible in memory. This significantly extends the receptive field of existing vision models with a marginal increase in memory cost and the number of parameters.
@@ 5 Experiments
We focus on iNaturalist 2018 [34] for classification, xView3-SAR [25] for segmentation, and MS-COCO [20] for detection. Since iNaturalist 2018 is a massive dataset, we focus on the Reptilia super-class, the most challenging subset available in the benchmark [34]. xView3-SAR is a difficult segmentation dataset for two reasons: it is comprised of extremely large images that are 29,400$\times$24,400 pixels large on average, and the objects in the dataset are heavily influenced by their non-local surroundings. Lastly, MS-COCO is used for evaluation of detection and to serve as a common baseline across prior work.
@@ 5.1 Classification
Table 2 contains results for variants of Swin, Hiera, and their xT variants on iNaturalist-Reptilia. Particularly, we demonstrate xT’s results as a function of region/input size and type of context encoder. xT outperforms their comparable baselines by significant margins.
@@ 5.2 Segmentation on xView3-SAR
We test Swin-Tiny, Swin-Small, and Swin-Base, and their xT-variants using Transformer-XL as a context encoder with $N=4$ layers. The models are trained over encoder input sizes of $512$ and $1024$, which xT effectively boosts to 4096. At inference time, we take overlapping crops of the same size across the image and combine our predictions with some post-processing. We sweep our models over learning rates of $\{3\times 10^{-5},10^{-4},3\times 10^{-4},10^{-3},3\times 10^{-3}\}$ using AdamW (with the same hyperparameters as in Section 5.1) and report the validation numbers in Table 3.
While prior works tackle xView3 using smaller, lighter models like CNNs, ours opts to adapt transformers for the task, which is a relatively unexplored space for the type of images and data distributions present. Therefore, larger models are not always better, as observed in Table 3. However, xT always outperforms the corresponding non-context model, especially on the overall $F_{1}$ detection score, signaling our models broad ability to improve models on large context tasks. We do this while only introducing $5-10\%$ extra parameters, and thanks to our nested tokenization, the memory we utilize per region is much lower as we see multiple regions at once per pass.
@@
[…]
Swin-B-DINO 30.0 36.1 256
Swin-B-DINO 32.5 39.6 512
Swin-B-DINO ⟨xT⟩ Hyper 49.8 68.5 1333/256
Swin-B-DINO ⟨xT⟩ Hyper 50.9 70.6 1333/512
Table 4: Comparison with Swin-B on the COCO 2017 object detection task. We report mAP and $\text{mAP}_{\text{Large}}$ on val2017.
@@ 5.3 Object Detection
We use SwinV2 [22] as the backbone and adopt DINO [41] for the detection head. We follow the standard COCO 1x training schedule and train on MS-COCO for 12 epochs with the AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a learning rate of $1\times 10^{-3}$. The learning rate for the SwinV2 backbone is scaled by a factor of 0.1. The input is first padded to the nearest multiple of chip size first, and then chipped in the same way as Sec. 5.1. Each region is then processed by the region encoder independently. The outputs from Swin encoder are concatenated to form a global feature map of the entire image, which is then processed by the context encoder. The resulting sequence is passed to the DINO detection head.
When utilizing the commonly adopted windowing approach, large objects that span multiple windows are split apart. This is a challenging task for state-of-the-art vision models to handle. Following, we report mAP and $\text{mAP}_{\text{Large}}$ for our detection experiments on val2017 subset. Comparing with plainly using Swin, xT  achieves strong performance when compared with baselines.
@@ Acknowledgements
Trevor Darrell’s group was supported in part by funding from the Department of Defense as well as BAIR’s industrial alliance programs. Ritwik Gupta is supported by the National Science Foundation under Grant No. DGE-2125913 and funding from the UC Berkeley AI Policy Hub. We are grateful to Bryan Comstock for his timely technical support. Thank you to Suzie Petryk, Mark Lowell, and William Yang for reviewing the transcript.

### v2

@@ Abstract
Modern computer vision pipelines handle large images in one of two sub-optimal ways: down-sampling or cropping. These two methods incur significant losses in the amount of information and context present in an image. There are many downstream applications in which global context matters as much as high frequency details, such as in real-world satellite imagery; in such cases researchers have to make the uncomfortable choice of which information to discard. We introduce xT, a simple framework for vision transformers which effectively aggregates global context with local details and can model large images end-to-end on contemporary GPUs. We select a set of benchmark datasets across classic vision tasks which accurately reflect a vision model’s ability to understand truly large images and incorporate fine details over large scales and assess our method’s improvement on them. xT is a streaming, two-stage architecture that adapts existing vision backbones and long sequence language models to effectively model large images without quadratic memory growth. We are able to increase accuracy by up to 8.6% on challenging classification tasks and $F_{1}$ score by 11.6 on context-dependent segmentation on images as large as 29,000 x 29,000 pixels.
@@ 1 Introduction
As camera technology has advanced, images have been getting increasingly larger over the past decade. Images captured by sensors on smartphones now capture images at 4K resolution (roughly 8.3M pixels) while professional DSLR cameras capture images at 8K resolution. Elsewhere, sensors on satellites and microscopes capture images with over a billion pixels.
@@
We introduce xT, a streaming, two-stage framework by which myopic vision backbones can effectively integrate local and global context over large images without incurring quadratic memory growth. In particular, we tackle both issues of increasing GPU memory utilization and the integration of context across very large images. We achieve this by introducing token hierarchies to state-of-the-art vision backbones [22, 31] and imbuing the resulting local features with global context through the use of long-sequence models, such as Transformer-XL and Mamba [4, 10], obtained from the field of natural language processing. xT matches and beats the performance of competitive large image architectures on multiple downstream tasks that require large visual contexts such as segmentation, detection, and classification. We demonstrate results on a variety of downstream tasks and achieve up to an 8.6% gain in accuracy on classification tasks and an 11.6 increase in $F_{1}$ score on context-dependent segmentation.
@@ 2 Related Works
We evaluate on real-world datasets such as xView3-SAR [26], a dataset where the average image is $29,400\times 24,400$ pixels large, and iNaturalist 2018 [35], where the entire $800\times 800$ image is utilized at once for evaluation. We also evaluate on the standard Cityscapes dataset for direct comparison on a common vision benchmark.
@@
xT is a streaming, two-stage architecture in which a powerful vision backbones extracts features from regions of a large image in batches. These features are streamed to a shallow context encoder which integrates global context across local features, effectively increasing the receptive field of the vision backbone across the entire image while staying within memory and parameter limits.
@@ 4 Methodology
Our goal for xT is to demonstrate a simple framework for allowing existing methods to process large images in a memory efficient and context-preserving manner. We achieve this through a streaming, two-stage architecture. First, images are tokenized in a multi-stage hierarchy (Section 4.1) into regions. These regions are encoded by a powerful, but myopic, vision encoder in batches (at worst case, in serial) (Section 4.2), resulting in a set of local features for the entire image. A lightweight context encoder then integrates information across all local features, culminating in a feature vector which contains contextualized global semantics over the entire image (Section 4.3). This feature vector is then used for task-specific decoding. Our overall pipeline is illustrated in Figure 2.
@@ 4.2 Region Encoder
The region encoder is any vision model which has been previously trained on small images small images $H\times W$, usually $224\times 224$ or $256\times 256$. The region encoder independently generates feature maps for each region $R^{i}_{1...P}$. In our experiments, we utilize vision transformers which output a shorter sequence length than which is input to them. These sequence lengths are less than the equivalent length produced by isotropic ViTs [6]. In this setup, we are able to effectively handle images with an increased number of regions, as our sequence length is reduced by 4×4$\times$4 × or greater. However, we also demonstrate the effectiveness of xT with a CNN-based encoder in the Appendix.
We stream regions through the region encoder in batches when GPU memory allows. However, either when the image is too large such that all of its constituent regions cannot fit into GPU memory, or when the regions themselves are too large, we process the image sequentially. The features generated from this process contain information limited to each region and are concatenated in row-major order to form a sequence.
@@ 4.3 Context Encoder
The context encoder is a lightweight sequence-to-sequence model that is able to effectively attend to long sequences. In xT, the context encoder plays the key role of disaggregating global context across local features extracted using the region encoder. Critically, we constrain the context encoder to have a near-linear memory and time cost. This design allows xT to “see” many other regions of the large image which otherwise would not be feasible with the naive usage of a vision model. xT’s usage of a context encoder significantly extends the receptive field of existing vision models with a marginal increase in memory cost and the number of parameters.
@@
Vision models used as-is are able to have an effective receptive field across the entire image but have zero reception on large images as they simply run out of memory. We aim to demonstrate the effective receptive field of xT’s two-stage architecture. In summary, the receptive field of xT is global but with minor aliasing artifacts due to an effective convolution over the image with the region encoder.
@@ 5 Experiments
We focus on iNaturalist 2018 [35] for classification, xView3-SAR [26] for segmentation, and Cityscapes [3] for detection. Since iNaturalist 2018 is a massive dataset, we focus on the Reptilia super-class, the most challenging subset available in the benchmark [35]. xView3-SAR is a difficult segmentation dataset for two reasons: it is comprised of extremely large images that are 29,400$\times$24,400 pixels large on average, and the objects in the dataset are heavily influenced by their non-local surroundings. Lastly, Cityscapes is used for evaluation of detection and to serve as a common baseline across prior work.
@@ 5.1 Classification
Table 2 contains results for variants of Swin, Hiera, and their xT variants on iNaturalist-Reptilia. Particularly, we demonstrate xT’s results as a function of region/input size and type of context encoder. xT outperforms their comparable baselines by up to 8.6% in top-1 accuracy.
@@ 5.2 Segmentation on xView3-SAR
We test Swin-T, Swin-S, and Swin-B, and their xT-variants using Transformer-XL as a context encoder with $N=4$ layers. The models are trained over encoder input sizes of $512$ and $1024$, which xT effectively boosts to 4096. At inference time, we take overlapping crops of the same size across the image and combine our predictions with post-processing. We sweep our models over learning rates of $\{3\times 10^{-5},10^{-4},3\times 10^{-4},10^{-3},3\times 10^{-3}\}$ using AdamW (with the same hyperparameters as in Section 5.1) and report the validation numbers in Table 3.
While prior works tackle xView3 using smaller models like CNNs, ours opts to adapt transformers for the task, which is an under-unexplored space for the type of images present. Therefore, larger models are not always better, as observed in Table 3. However, xT always outperforms the corresponding non-context model, beating baselines by up to 13.4 points on the overall $F_{1}$ detection score and 6.0 points on the close-to-shore $F_{1}$ score, signaling xT’s ability to improve models on large context tasks. Due to our streaming, two-stage architecture with nested tokenization, the memory utilized per region is much lower as we see multiple regions at once per pass while only introducing $5-10\%$ more parameters, .
@@
Model mAP Input Size(s)
Swin-B-DetINO OOM 2048
Swin-B-DetINO ⟨xT⟩ XL 43.0 2048/512
Table 4: Comparison with Swin-B on the Cityscapes object detection task. Swin-B is unable to model Cityscapes images in their entirety within the memory of an 80GB A100.
@@ 5.3 Object Detection
We use SwinV2 [23] as the backbone and adopt DetINO [42] for the detection head. We follow the standard COCO [21] 1x training schedule and train on Cityscapes for 12 epochs with the AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and a learning rate of $1\times 10^{-3}$. The learning rate for the SwinV2 backbone is scaled by a factor of 0.1. The input is first padded to the nearest multiple of chip size first, and then chipped in the same way as Sec. 5.1. Each region is then processed by the region encoder independently. The outputs from Swin encoder are concatenated to form a global feature map of the entire image, which is then processed by the context encoder. The resulting sequence is passed to the DetINO detection head.
We report mAP for our detection experiments. These experiments are relatively brief—naïve usage of the Swin backbone results in out-of-memory errors for the large 1,024 x 2,048 images found in Cityscapes. Conversely, xT is able to model the images with ease.
@@ 7.3 Memory Growth
7.4 Locality of Features
Many vision backbones have adjustable receptive fields for the integration of local information. Larger receptive fields have memory and throughput implications, however. We ablate the Swin transformer for various region and window sizes and compare the impact on accuracy, throughput, and memory in Table 6 on the iNaturalist classification task.
We set the window size used by Swin appropriately as a function of region size. That is, an region size of 256 corresponds to a window size of 16, region size of 512 to a window size of 32, and so on. Swin-B, without the usage of xT, can achieve marginally higher accuracy at small image sizes. However, this trend is reversed when working with larger images. Of importance is the ability for xT to retain high throughput (measured as time per epoch in minutes) and low memory usage (memory per region in GB) while achieving equivalent accuracy downstream.
Model Acc. Size(s) W. Size Th. Mem.
Swin-B 67.02 512 32 29.53 6.02
Swin-B ⟨xT⟩ XL 65.49 512/256 16 12.77 0.49
Swin-B 67.37 1024 64 242.92 25.1
Swin-B ⟨xT⟩ XL 68.19 1024/256 16 64.7 2.03
Table 6: Comparison of top-1 accuracy (Acc.), time per epoch (Th.), and memory per region (Mem.) as a function of region size(s) and window sizes. Swin-B and our method have approximately equivalent accuracies, but xT achieves much more desirable throughput and memory utilization.
@@ 8 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
Trevor Darrell’s group was supported in part by funding from the Department of Defense as well as BAIR’s industrial alliance programs. This work was additionally supported by the DoD’s High Performance Computing Modernization Program. Ritwik Gupta is supported by the National Science Foundation under Grant No. DGE-2125913 and funding from the UC Berkeley AI Policy Hub. We are grateful to Bryan Comstock for his timely technical support even when the authors were asking a lot of questions. Thank you to Suzie Petryk, Mark Lowell, and William Yang for reviewing the transcript.